# ThinkRec: Thinking-based recommendation via LLM

### v1

@@ Abstract
Recent advances in large language models (LLMs) have enabled more semantic-aware recommendations through natural language generation. Existing LLM for recommendation (LLM4Rec) methods mostly operate in a System 1-like manner, relying on superficial features to match similar items based on click history, rather than reasoning through deeper behavioral logic. This often leads to superficial and erroneous recommendations. Motivated by this, we propose ThinkRec, a thinking-based framework that shifts LLM4Rec from System 1 to System 2 (rational system). Technically, ThinkRec introduces a thinking activation mechanism that augments item metadata with keyword summarization and injects synthetic reasoning traces, guiding the model to form interpretable reasoning chains that consist of analyzing interaction histories, identifying user preferences, and making decisions based on target items. On top of this, we propose an instance-wise expert fusion mechanism to reduce the reasoning difficulty. By dynamically assigning weights to expert models based on users’ latent features, ThinkRec adapts its reasoning path to individual users, thereby enhancing precision and personalization. Extensive experiments on real-world datasets demonstrate that ThinkRec significantly improves the accuracy and interpretability of recommendations  [^1] Our implementations are available in anonymous Github: https://anonymous.4open.science/r/ThinkRec_LLM.
{CJK}
UTF8gbsn
ThinkRec: Thinking-based recommendation via LLM
Qihang Yu Zhejiang University yuqihang@zju.edu.cn                        Kairui Fu Zhejiang University fukairui.fkr@zju.edu.cn                        Shengyu Zhang Zhejiang University sy_zhang@zju.edu.cn
Zheqi Lv Zhejiang University zheqilv@zju.edu.cn                        Fan Wu Shanghai Jiao Tong University fwu@cs.sjtu.edu.cn                        Fei Wu Zhejiang University wufei@zju.edu.cn
Figure 1: (a) shows the composition of user interaction data. (b) and (c) illustrate previous LLM-based recommendations and our ThinkRec, respectively. (d) compares ThinkRec with baselines in three real-world datasets.
@@ 1 Introduction
Recommendation systems are indispensable in modern digital platforms, enabling users to navigate vast content efficiently Glenski et al. (2021); Ma et al. (2023); Iana et al. (2024). Traditional sequential recommendation methods rely on implicit modeling of user interaction histories, cannot model context and incorporate broader knowledge, which restricts their reasoning and generalization ability. Recent advances in LLMs offer strong contextual comprehension and extensive world knowledge to improve recommendation systems.
Prior solution paradigms of LLM4Rec can be seperated into three categories: (1) item scoring Liu et al. (2024); Zhang et al. (2025), where LLMs answer binary preference questions given user and item context; (2) item generation Li et al. (2023); Bao et al. (2024), which maps natural language prompts to item IDs through aligned embedding spaces; and (3) hybrid modeling Geng et al. (2023); Zhang et al. (2024b), where a single LLM performs multiple tasks such as pointwise prediction, pairwise ranking, or listwise generation. Although these LLM4Rec methods differ in output formats and representation learning strategies, they fundamentally resemble System 1—the intuitive system—in cognitive science Kahneman (2011). They tend to match similar items based on click history while overlooking the deeper behavioral logic. This limitation becomes evident in cases such as the one shown in Figure 1(b), where the user’s behavior over time is: dislikes “Dune”, likes “The Three-Body Problem”, and likes “Foundation” (all three are science fiction). Methods that rely on System 1-like intuition tend to infer that the user would also like “Hyperion” (science fiction) simply because it belongs to the same genre. In reality, the user might dislike philosophical or metaphysical themes in fiction, which are prominent in *Hyperion*, making it an unsuitable recommendation. Clearly, if we can leverage the vast knowledge encoded in LLMs and fully activate their reasoning capabilities for user behavior modeling, recommendation performance can be significantly improved.
@@ 2.1 LLM-based Recommendation
With the rapid progress of LLMs, there has been growing interest in adapting them to recommendation tasks using both textual and structured data. Early works Bao et al. (2023); Ho et al. (2023); Lin et al. (2024a); Zhang et al. (2025) rely on prompt-based scoring by converting recommendations into binary question-answering, while others Ji et al. (2023a); Li et al. (2023); Lin et al. (2024b) align language space with item embeddings to directly generate item IDs. More advanced approaches, such as P5 Geng et al. (2022) and InstructRec Zhang et al. (2024b) unify multiple sub-tasks—like rating prediction, pairwise comparison, or ranking, into a single language modeling framework. These methods demonstrate LLMs’ flexibility in expressing recommendation semantics, yet most focus on adapting output formats, treating LLMs as static scorers or selectors with limited interpretability.
Beyond static prompting, recent efforts shift toward viewing LLMs as interactive agents capable of multi-step reasoning and personalized decision-making. For example, RecMind Wang et al. (2024a) and MACRec Wang et al. (2024b) introduce modularized reasoning structures, enabling planning-based recommendation. Others explore integrating LLMs into various stages of the pipeline—for instance, generating user or item representations Zhang et al. (2024c) or acting as high-level decision controllers Park and Zhang (2025). Meanwhile, instruction tuning Zhang et al. (2024b), chain-of-thought prompting Liu et al. (2025) have been proposed to enhance LLMs’ semantic understanding. Despite these developments, explicit reasoning supervision and fine-grained user modeling remain underexplored. ThinkRec addresses these gaps by integrating reasoning-augmented training and personalized expert fusion, enabling interpretable, user-specific recommendations grounded in structured reasoning.
@@ 2.2 Reasoning Model
Recent advances have transformed LLMs from passive token predictors into structured reasoning agents by incorporating process-level supervision. Techniques such as Chain-of-Thought prompting Wei et al. (2022), ReAct Yao et al. (2022), and Tree-of-Thoughts Yao et al. (2023) enable models to generate intermediate reasoning steps, improving interpretability and multistep inference quality. To further reinforce reasoning capabilities, recent works have introduced process reward models Zhang et al. (2024a) and self-improvement pipelines using techniques like Monte Carlo Tree Search Luo et al. (2024) and reasoning distillation Xu et al. (2023). These trends define the emerging paradigm of Large Reasoning Models, which prioritize explicit, verifiable thinking processes over direct output. ThinkRec draws inspiration from this direction by introducing reasoning-augmented training for recommendation, aligning to embed structured reasoning into decision making.
@@ 3.1 Preliminary
Integrating Collaborative Embeddings into LLMs. The sequential recommenders, such as MF Koren et al. (2009) and LightGCN He et al. (2020), are used for encoding collaborative information, which can be integrated into LLM embeddings. Formally, for each sample $(u,t,i_{id},l)$:
@@ 3.2.1 Reasoning data construction
To support LLMs for more effective reasoning, we need more textual information in addition to item titles. Given that the raw metadata often contains excessive or noisy content, directly using long-form text may hinder the understanding and inference of LLMs. To address this, we leverage a pre-trained summarization model Wei et al. (2023) to extract key concepts from the item metadata, and retain up to 10 representative keywords to enhance the semantic representation of each item. To model more comprehensive preference information, not just liking information, we add low-scoring interactions to the history sequence as well. Then we augment each item with title $i_{t}$, label $i_{l}\in\{yes,no\}$, feature, and keywords $i_{k}$ as follows:
@@
Since existing recommendation datasets lack explicit reasoning traces, we synthesize a subset of samples with explainable rationale by leveraging a powerful reasoning model QwQ [^2] https://qwenlm.github.io/blog/qwq-32b/. Specifically, we sample a few thousand instances from the training data and construct textual prompts according to Appendix C.2. We then repeatedly query the reasoning model until a correct prediction is produced and record the latest explanation content as the reason for interaction as in Figure 2(b).
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on three datasets: ML1M refers to MovieLens-1M [^3] https://grouplens.org/datasets/movielens/1m/. Yelp refers to the Yelp Open dataset [^4] https://business.yelp.com/data/resources/open-dataset/. Book refers to the “book” subset of the Amazon Product Review dataset [^5] https://nijianmo.github.io/amazon/index.html. The data processing details are in the Appendix A.1, and the statistical information on the processed datasets is in Table 1. $\sigma$, $\mu$ refer to the standard deviation and mean of user interactions.
Evaluation Metrics. We employ four widely used recommendation metrics: AUC, UAUC, Normalized Discounted Cumulative Gain (NDCG), and Mean Average Precision (MAP). And we employ METEOR Banerjee and Lavie (2005) and BLEURT Sellam et al. (2020) to measure the generated reasons. METEOR incorporates synonym matching and word order. BLEURT is fine-tuned on human judgments to directly predict text quality.
Baselines. To fully evaluate the proposed method ThinkRec, we compare it with traditional collaborative methods MF Koren et al. (2009), LightGCN He et al. (2020), SASRec Kang and McAuley (2018), and LLM-based methods Prompt4NR Zhang and Wang (2023), TALLRec Bao et al. (2023), CoLLM Zhang et al. (2025). We extend the above LLM-based method to LLM Llama3-8B666huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for a fair comparison and tune the LLM with LoRA to manage computational costs. The complete introduction is in Appendix A.3.
Implementation Details We implement all the compared methods using PyTorch 2.5. We adopt BCE when not otherwise specified. LLM-based methods are optimized using AdamW, while other models use Adam. For hyperparameters such as learning rate, embedding size, and weight decay of traditional methods, we conduct grid search over commonly used ranges, and follow the original papers for other baseline-specific settings. For LLM-based methods, we adopt CoLLM’s configuration for optimizer, sequence truncation and LoRA settings. We set the sampling ratio of reasoning data to 0.2, and full hyperparameter settings and search ranges are provided in the Appendix A.2.
@@ 4.2 Overall Performance (RQ1)
We compare our ThinkRec method with baseline methods on the accuracy of recommendations and the quality of generated reasons.
@@ 4.2.1 Accuracy of Recommendation
Table 3: Quality evaluation of generated reasons. "M" refers to "METEOR" and "B" refers to "BLEURT".
4.2.2 Quality of Generated Reasons
Table 3 summarizes the quality of the reasons generated by the LLM-based recommenders. Compared with the generated reasons from QwQ using the METEOR and BLEURT metrics, the reasons generated by our method significantly outperform those of the three LLM-based baselines. Our method achieves an average relative improvement of 56.54% on METEOR and 23.35% on BLEURT across all datasets, suggesting better fluency, coherence, and semantic relevance. These results validate the effectiveness of our thinking activation mechanism, which explicitly aligns recommendations with user-centric reasoning via joint training on reasoning-augmented samples. The improvement in both syntactic and learned semantic metrics confirms that ThinkRec not only provides accurate recommendations but also produces more coherent, grounded, and human-aligned explanations, a crucial step toward reasonable and trustworthy LLM-based recommendations.
@@ 4.2.2 Quality of Generated Reasons
4.3 In-depth Analysis
4.3.1 Ablation Studies (RQ2)
To evaluate the importance of explicit reasoning in recommendation, we ablate the "thinking" component of ThinkRec (w/o think), which disables reasoning supervision. As shown in Table 4, this leads to significant performance degradation across all datasets. For example, UAUC on the Book dataset drops from 0.5705 to 0.4692. Interestingly, this is even lower than the variant where both thinking and expert mechanisms are removed (w/o both). Without thinking, the recommendation task is effectively reduced to binary classification, where enabling multi-expert modeling may overfit shallow interactions, thus compromising generalization.
@@ 4.3.2 Study on the Fusion of Experts (RQ3)
Analysis of the number of experts. As shown in Figure 3, the number of expert groups increases from 1 to 4, the model exhibits a characteristic ’rise-then-fall’ performance trend, revealing the trade-off between personalization capacity and generalization. In the early stages, fine-tuning LoRA modules within user groups significantly enhances the model’s ability to capture diverse preferences, resulting in notable gains in user-level and Top-N ranking metrics such as UAUC, NDCG@5, and MAP. However, with further partitioning, each subgroup receives fewer training samples, making the model prone to overfitting, thereby degrading ranking performance. Notably, the AUC metric consistently decreases with more experts, reflecting the deterioration of the consistency of global representation with expert specialization, validating the inherent tension between "global consistency" and "local specificity" in recommender systems. These results indicate that more experts do not necessarily equate to better performance; instead, the optimal group number should be dynamically adjusted according to user behavior diversity, and the frequency of interactions.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have enabled more semantic-aware recommendations through natural language generation. Existing LLM for recommendation (LLM4Rec) methods mostly operate in a System 1-like manner, relying on superficial features to match similar items based on click history, rather than reasoning through deeper behavioral logic. This often leads to superficial and erroneous recommendations. Motivated by this, we propose ThinkRec, a thinking-based framework that shifts LLM4Rec from System 1 to System 2 (rational system). Technically, ThinkRec introduces a thinking activation mechanism that augments item metadata with keyword summarization and injects synthetic reasoning traces, guiding the model to form interpretable reasoning chains that consist of analyzing interaction histories, identifying user preferences, and making decisions based on target items. On top of this, we propose an instance-wise expert fusion mechanism to reduce the reasoning difficulty. By dynamically assigning weights to expert models based on users’ latent features, ThinkRec adapts its reasoning path to individual users, thereby enhancing precision and personalization. Extensive experiments on real-world datasets demonstrate that ThinkRec significantly improves the accuracy and interpretability of recommendations .
@@ 1 Introduction
Recommendation systems are indispensable in modern digital platforms, enabling users to navigate vast content efficiently [1, 2, 3]. Traditional sequential recommendation methods rely on implicit modeling of user interaction histories, cannot model context and incorporate broader knowledge, which restricts their reasoning and generalization ability. Recent advances in LLMs offer strong contextual comprehension and extensive world knowledge to improve recommendation systems.
Prior solution paradigms of LLM4Rec can be seperated into three categories: (1) item scoring [4, 5], where LLMs answer binary preference questions given user and item context; (2) item generation [6, 7], which maps natural language prompts to item IDs through aligned embedding spaces; and (3) hybrid modeling [8, 9], where a single LLM performs multiple tasks such as pointwise prediction, pairwise ranking, or listwise generation. Although these LLM4Rec methods differ in output formats and representation learning strategies, they fundamentally resemble System 1—the intuitive system—in cognitive science [10]. They tend to match similar items based on click history while overlooking the deeper behavioral logic. This limitation becomes evident in cases such as the one shown in Figure 1(b), where the user’s behavior over time is: dislikes “Dune”, likes “The Three-Body Problem”, and likes “Foundation” (all three are science fiction). Methods that rely on System 1-like intuition tend to infer that the user would also like “Hyperion” (science fiction) simply because it belongs to the same genre. In reality, the user might dislike philosophical or metaphysical themes in fiction, which are prominent in *Hyperion*, making it an unsuitable recommendation. Clearly, if we can leverage the vast knowledge encoded in LLMs and fully activate their reasoning capabilities for user behavior modeling, recommendation performance can be significantly improved.
Figure 1: (a) shows the composition of user interaction data. (b) and (c) illustrate previous LLM-based recommendations and our ThinkRec, respectively. (d) compares ThinkRec with baselines.
@@ 2.1 LLM-based Recommendation
With the rapid progress of LLMs, there has been growing interest in adapting them to recommendation tasks using both textual and structured data. Early works [11, 12, 13, 5] rely on prompt-based scoring by converting recommendations into binary question-answering, while others [14, 6, 15] align language space with item embeddings to directly generate item IDs. More advanced approaches, such as P5 [16] and InstructRec [9] unify multiple sub-tasks—like rating prediction, pairwise comparison, or ranking, into a single language modeling framework. These methods demonstrate LLMs’ flexibility in expressing recommendation semantics, yet most focus on adapting output formats, treating LLMs as static scorers or selectors with limited interpretability.
Beyond static prompting, recent efforts shift toward viewing LLMs as interactive agents capable of multi-step reasoning and personalized decision-making. For example, RecMind [17] and MACRec [18] introduce modularized reasoning structures, enabling planning-based recommendation. Others explore integrating LLMs into various stages of the pipeline—for instance, generating user or item representations [19] or acting as high-level decision controllers [20]. Meanwhile, instruction tuning [9], chain-of-thought prompting [21] have been proposed to enhance LLMs’ semantic understanding. Despite these developments, explicit reasoning supervision and fine-grained user modeling remain underexplored. ThinkRec addresses these gaps by integrating reasoning-augmented training and dynamic expert fusion, enabling interpretable, user-specific recommendations grounded in reasoning.
@@ 2.2 Reasoning Model
Recent advances have transformed LLMs from passive token predictors into structured reasoning agents by incorporating process-level supervision. Techniques such as Chain-of-Thought prompting [22], ReAct [23], and Tree-of-Thoughts [24] enable models to generate intermediate reasoning steps, improving interpretability and multistep inference quality. To further reinforce reasoning capabilities, recent works have introduced process reward models [25] and self-improvement pipelines using techniques like Monte Carlo Tree Search [26] and reasoning distillation [27]. These trends define the emerging paradigm of Large Reasoning Models, which prioritize explicit, verifiable thinking processes over direct output. ThinkRec draws inspiration from this direction by introducing reasoning-augmented training for recommendation, aligning to embed structured reasoning into decision making.
@@ 3.1 Preliminary
Integrating Collaborative Embeddings into LLMs. The sequential recommenders, such as MF [28] and LightGCN [29], are used for encoding collaborative information, which can be integrated into LLM embeddings. Formally, for each sample $(u,t,i_{id},l)$:
@@ 3.2.1 Reasoning data construction
To support LLMs for more effective reasoning, we need more textual information in addition to item titles. Given that the raw metadata often contains excessive or noisy content, directly using long-form text may hinder the understanding and inference of LLMs. To address this, we leverage a pre-trained summarization model [30] to extract key concepts from the item metadata, and retain up to 10 representative keywords to enhance the semantic representation of each item. To model more comprehensive preference information, not just liking information, we add low-scoring interactions to the history sequence as well. Then we augment each item with title $i_{t}$, label $i_{l}\in\{yes,no\}$, feature, and keywords $i_{k}$ as follows:
@@
Since existing recommendation datasets lack explicit reasoning traces, we synthesize a subset of samples with explainable rationale by leveraging a powerful reasoning model QwQ [^1] https://qwenlm.github.io/blog/qwq-32b/. Specifically, we sample a few thousand instances from the training data and construct textual prompts according to Appendix B.3. We then repeatedly query the reasoning model until a correct prediction is produced and record the latest explanation content as the reason for interaction as in Figure 2(b).
@@ 4.1 Experimental Setup
Datasets.
@@
We conduct experiments on three datasets: ML1M refers to MovieLens-1M [^2] https://grouplens.org/datasets/movielens/1m/. Yelp refers to the Yelp Open dataset [^3] https://business.yelp.com/data/resources/open-dataset/. Book refers to the “book” subset of the Amazon Product Review dataset [^4] https://nijianmo.github.io/amazon/index.html. ML1M contains user ratings on movies, collected between 2000 and 2003, with ratings on a scale of 1 to 5. We convert these ratings into binary labels using a threshold of 3. Yelp includes user reviews, ratings for businesses such as restaurants and retail shops, as well as textual information about the businesses. We convert these ratings into binary labels using a threshold of 3. Book compiles user reviews of books from Amazon, collected between 1996 and 2018, with review scores ranging from 1 to 5. We transform these review scores into binary labels using a threshold of 4. Ratings greater than the threshold are labeled as “positive” (y = 1), while the rest are labeled as “negative” (y = 0).
To better simulate real-world recommendation scenarios and prevent data leakage [31], we split the dataset into training, validation, and testing sets based on the interaction timestamp. Specifically, for Amazon-Book, given its large scale, we just preserve interactions from the year 2017, allocating the first 10 months for training, and the remaining 2 months for validation and testing. As for Yelp, we preserve interactions from 2010 to 2022, allocating the first 10 years for training, and the remaining 2 years for validation and testing. Given the sparse nature of Book and Yelp, we filtered out users and items with fewer than 20 interactions. As for ML1M, we preserve the interactions from the most recent twenty months, using the first 10 months for training, the remaining 10 months for validation and testing. The statistical information on the processed datasets is in Table 1. $\sigma$, $\mu$ refer to the standard deviation and mean of user interactions.
Evaluation Metrics. We employ four widely used recommendation metrics: AUC, UAUC, Normalized Discounted Cumulative Gain (NDCG), and Mean Average Precision (MAP). And we employ METEOR [32] and BLEURT [33] to measure the generated reasons. METEOR incorporates synonym matching and word order. BLEURT is fine-tuned on human judgments to directly predict text quality.
Baselines. To fully evaluate the proposed method ThinkRec, we compare it with traditional collaborative methods, and LLM-based methods. We extend LLM-based methods to LLM Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for a fair comparison and tune the LLM with LoRA to manage computational costs.
  • MF [28] is Matrix Factorization, a latent factor-based collaborative filtering method.
  • LightGCN [29] A representative graph-based collaborative filtering method, which uses a simplified graph convolutional neural network to enhance the modeling of user interest.
  • SASRec [34] A representative sequential recommendation method, which uses self-attention to encode sequential patterns to model user interest.
  • Prompt4NR [35] It uses both fixed and soft prompts to utilize traditional Language Models for recommendation.
  • TALLRec [11] This is a state-of-the-art LLMRec method that aligns LLM with recommendations through aligns LLM with recommendations through instruction tuning.
  • CoLLM [5] It effectively integrates collaborative information into LLMs by harnessing the capability of external traditional models to capture the information.
Implementation Details We implement all the compared methods using PyTorch 2.5. We adopt BCE when not otherwise specified. LLM-based methods are optimized using AdamW, while other models use Adam. For hyperparameters of traditional methods, we explore the learning rate within the range of [1e-2, 1e-3, 1e-4], the embedding size within the range of [64, 128, 256], and weight decay within the range of [1e-2, 1e-3, . . . , 1e-7]. As for LLM-based methods, we set the learning rate to 1e-4, and weight decay to 1e-3 to align with CoLLM. For SASRec, we establish the maximum length of historical interaction sequences to 25. We adopt TALLRec’s practice of setting the maximum sequence length to 10 for all other methods. Regarding other specific parameters of the baseline models, we adhere to the configurations outlined in their original papers. For the LoRA module, we follow the same configuration as CoLLM, setting $r,alpha,dropout,target\ modules$ to 8, 16, 0.05, and “[q proj, v proj]”, respectively. We set the sample rate of reasoning data and recommendation data to 0.2 and 0.8, and the weight of each loss function as follows: $\alpha=0.1,\beta=0.9,\eta=0.9,\gamma=0.1$. We set the temperature coefficient $\tau$ to $0.1$. And we set the sampling ratio of reasoning data to 0.2.
@@ 4.2 Overall Performance (RQ1)
In this section, we compare our ThinkRec method with baseline methods on the accuracy of recommendations and the quality of generated reasons.
@@ 4.2.2 Quality of Generated Reasons
Table 3: Evaluation quality of generated reasons. "M" refers to "METEOR" and "B" refers to "BLEURT".
Table 3 summarizes the quality of the reasons generated by the LLM-based recommenders. Compared with the generated reasons from QwQ using the METEOR and BLEURT metrics, the reasons generated by our method significantly outperform those of the three LLM-based baselines. Our method achieves an average relative improvement of 56.54% on METEOR and 23.35% on BLEURT across all datasets, suggesting better fluency, coherence, and semantic relevance.
These results validate the effectiveness of our thinking activation mechanism, which explicitly aligns recommendations with user-centric reasoning via joint training on reasoning-augmented samples. The improvement in both syntactic and learned semantic metrics confirms that ThinkRec not only provides accurate recommendations but also produces more coherent, grounded, and human-aligned explanations, a crucial step toward reasonable and trustworthy LLM-based recommendations.
4.3 In-depth Analysis
4.3.1 Ablation Studies (RQ2)
@@
To evaluate the importance of explicit reasoning in recommendation, we ablate the "thinking" component of ThinkRec (w/o think), which disables reasoning supervision.
As shown in Table 4, this leads to significant performance degradation across all datasets. For example, UAUC on the Book dataset drops from 0.5705 to 0.4692. Interestingly, this is even lower than the variant where both thinking and expert mechanisms are removed (w/o both). Without thinking, the recommendation task is effectively reduced to binary classification, where enabling multi-expert modeling may overfit shallow interactions, thus compromising generalization.
@@ 4.3.2 Study on the Fusion of Experts (RQ3)
Analysis of the number of experts. As shown in Figure 3, the number of expert groups increases from 1 to 4, the model exhibits a characteristic ’rise-then-fall’ performance trend, revealing the trade-off between personalization capacity and generalization. In the early stages, fine-tuning LoRA modules within user groups significantly enhances the model’s ability to capture diverse preferences, resulting in notable gains in user-level and Top-N ranking metrics such as UAUC, NDCG@5, and MAP. However, with further partitioning, each subgroup receives fewer training samples, making the model prone to overfitting, thereby degrading ranking performance. Notably, the AUC metric consistently decreases with more experts, reflecting the deterioration of the consistency of global representation with expert specialization, validating the inherent tension between "global consistency" and "local specificity" in recommender systems. These results indicate that more experts do not necessarily equate to better performance; instead, the optimal group number should be adjusted according to user behavior diversity and the frequency of interactions.